I'm trying to figure out how to work a problem like this:
Evaluate each expression if a= 1 3/8, b= 4 5/7, c= 3 3/10
problem is
B divided A

Answer by stanbon(75887)

Evaluate each expression if a= 1 3/8, b= 4 5/7, c= 3 3/10
problem is
B divided A
----------
b/a = [33/7]/[11/8]
Invert the denominator and multiply to get:
=[33/7]
=[3/7]
=24/7
=3 3/7
Cheers,
Stan H.